• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Humensky

    Log entry time 08:16:59 on June26,2004

    Entry number 126938

    keyword=OWL Shift Summary

    00:00 IHWP -> OUT; resume production running with Run 2416
    01:00 Profile scan, full area -- run 2418 (2417 brief; wrong voltages 
          on profile detectors).  Scanner potentiometers had readout
          problems, so we decide to defer profile scan at half operating
          current.
    01:19 Resume production running
    02:10 Noticed Hall A IA loop was railing though charge asym still 
          zero.  Culprit was recent turn-on of Hall C beam (to 50uA).
    02:15 MCC called to say Hall A laser power is maxed out giving us 32uA
          because Hall C is getting 50uA (and draining cathode?).  With
          cathode lifetime short, we see current dropping steadily through
          remainder of shift.
    03:28 Raised lumi voltages to bring lumi signals up to ~40k at 30uA
          (run 2421).
    04:30 Inserted IHWP and updated RHWP, Pockels cell voltages, and Hall
          C IA voltages.
    05:00 Target IOC reboot.
    07:30 Noticed Left Septum He level had been dropping steadily for last ~2hrs.  Was at 51% when we caught it.  Called RC and then Tech on call.  Going into controlled access to reset He reservoir JT valve.
    
    HAPPEX DAQ run summary
    2416	production; removed IHWP prior to start.  IHWP OUT
    2417	profile scan; aborted after 3900 events (wrong prof det HV)
    2418	profile scan, full area.
    2419	production IHWP OUT
    2420	production IHWP OUT
    2421	production?  raising LUMI voltages during run but otherwise
    	normal.
    2422	production IHWP OUT
    2423	production; inserted IHWP prior to start.  IHWP IN
    2424	production IHWP IN
    2425	production IHWP IN; PITA fdbk at 45k events
    2426	production IHWP IN
    
    Compton DAQ run summary
    8516	IHWP OUT ~00:00 - 02:10
    8517	IHWP OUT  02:16 - 04:20 forgot to start left-right process and
    	when "end run" failed I never got an end-of-run window.
    8518	IHWP  IN  04:25 - 06:25 forgot to start left-right process
    8519	junk due to DAQ problem(?) -- after hitting "start run", 
    	window that opened crashed.
    8520	IHWP IN 06:55 - started left-right process.
    
    BCM Temperature summary
    Chan	1:50am		7:25am
    2	94.4		104.1
    3	94.3		103.4
    4	105.7		107.7
    5	106.1		108.0
    6	 70.2		 70.1